If n=150 and ˆpp^ (p-hat) =0.67, find the margin of error at a 95% confidence level

Give your answer to three decimals
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = 0.67

1 - \hat p = 1 - 0.67 = 0.33

Z\alpha/2 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 * (\sqrt((0.67 * 0.33) / 150)

Margin of error = E = 0.075